OpenStudy (anonymous):

@KingGeorge Prove that if m^2 | n^2, then m | n

OpenStudy (anonymous):

I tried doing k = (n/m)^2 but then I got stuck. Maybe I was approaching it wrong?

OpenStudy (kinggeorge):

It seems like that should work. You just have to note that \(k\) is a perfect square, and so when you take the square root, you get an integer.

OpenStudy (anonymous):

How do you know that it is a perfect square?

OpenStudy (kinggeorge):

Alternatively, let \(p\) be a prime dividing \(m\). Then \(p^2|n^2\). Thus, \(p|n^2\), so \(p\) must divide \(n\). That means that \(m|n\).

OpenStudy (kinggeorge):

And that was my mistake. We don't know offhand that it's a perfect square.

OpenStudy (anonymous):

Wait, yes, we do. Since m^2 | n^2, then m <= n, then our expression n/m must not be a fraction.

OpenStudy (anonymous):

Which means that (n/m)^2 must be a perfect square

OpenStudy (anonymous):

Since m^2 | n^2, then m <= n, then our expression n/m must be an integer

OpenStudy (kinggeorge):

Nice. So it looks like we have two ways to prove this now :)

OpenStudy (anonymous):

You don't see any faults in my logic do you?

OpenStudy (kinggeorge):

It looks fine to me. We know that \(n^2=m^2\cdot k\) for some integer \(k\), so \(k=n^2/m^2\). But using properties of squares, we get that \(k=(n/m)^2\), so it must be that \(n/m\) is an integer \(k_0\), which then implies that \(m|n\).